# Adaptive Additive Parameter Updates of Vision Transformers for Few-Shot Continual Learning

### v1

@@ Abstract
Integrating new class information without losing previously acquired knowledge remains a central challenge in artificial intelligence, often referred to as catastrophic forgetting. Few-shot class incremental learning (FSCIL) addresses this by first training a model on a robust dataset of base classes and then incrementally adapting it in successive sessions using only a few labeled examples per novel class. However, this approach is prone to overfitting on the limited new data, which can compromise overall performance and exacerbate forgetting. In this work, we propose a simple yet effective novel FSCIL framework that leverages a frozen Vision Transformer (ViT) backbone augmented with parameter-efficient additive updates. Our approach freezes the pre-trained ViT parameters and selectively injects trainable weights into the self-attention modules via an additive update mechanism. This design updates only a small subset of parameters to accommodate new classes without sacrificing the representations learned during the base session. By fine-tuning a limited number of parameters, our method preserves the generalizable features in the frozen ViT while reducing the risk of overfitting. Furthermore, as most parameters remain fixed, the model avoids overwriting previously learned knowledge when small novel data batches are introduced. Extensive experiments on benchmark datasets demonstrate that our approach yields state-of-the-art performance compared to baseline FSCIL methods.
@@ V-A Main Experimental Results
On CUB-200, our method improves the base accuracy ($S_{\mathrm{base}}$) by 3.25% and the final session accuracy ($S_{\mathrm{last}}$) by 9.18% over the best competing methods, while reducing the performance drop (PD) by 0.66%. These improvements show that our selective parameter updates not only preserve base knowledge but also better integrate new classes over incremental sessions.
On CIFAR-100, we achieve state-of-the-art accuracies during both the base and all incremental sessions. Although our PD (5.17%) is marginally higher than that of PriViLege (4.82%), our method avoids the extra complexity and parameter overhead associated with prompt tuning and word embeddings required by PriViLege.
For miniImageNet, our approach attains stellar performance across all sessions, setting a new benchmark PD of 1.29%, a new base session accuracy of 98.13% and final session accuracy of 96.84% when using the ViT backbone fine-tuned on ImageNet-1K. Interestingly, the less powerful backbone (with ImageNet-1K fine-tuning) outperforms its counterpart, likely due to better alignment with the miniImageNet distribution.
Overall, these results validate that our FSCIL framework, with its targeted additive updates and refined classifier weights, effectively maintains robust representations while accommodating new classes, leading to significant improvements over existing methods.

### v2

@@ Abstract
Integrating new class information without losing previously acquired knowledge remains a central challenge in artificial intelligence, often known as catastrophic forgetting. Few-shot class incremental learning (FSCIL) addresses this by first training a model on a robust set of base classes and then incrementally adapting it in successive sessions using few labeled examples per novel class. However, this approach is prone to overfitting on limited new data, which can compromise performance and exacerbate forgetting. In this work, we propose a simple yet effective FSCIL framework that leverages a frozen Vision Transformer (ViT) backbone augmented with parameter-efficient additive updates. Our approach freezes the pre-trained ViT parameters and selectively injects trainable weights into the self-attention modules via an additive update mechanism. This design updates a small subset of parameters to accommodate new classes without sacrificing representations learned during the base session. By fine-tuning only a few parameters, our method preserves the generalizable features in the frozen ViT while reducing overfitting risk. Furthermore, as most parameters remain fixed, the model avoids overwriting knowledge when novel data batches are introduced. Extensive experiments on benchmark datasets demonstrate that our approach yields state-of-the-art performance compared to baseline FSCIL methods. Our results confirm improvements in robustness and accuracy.
@@ V-A Main Experimental Results
On CUB-200, our method improves the base accuracy ($S_{\mathrm{base}}$) by 3.25% and the final session accuracy ($S_{\mathrm{last}}$) by 9.18% over the best competing methods, while reducing the performance drop (PD) by 0.66%. These improvements show that our selective parameter updates not only preserve base knowledge but also better integrate new classes over incremental sessions. On CIFAR-100, we achieve state-of-the-art accuracies during both the base and all incremental sessions. Although our PD (5.17%) is marginally higher than that of PriViLege (4.82%), our method avoids the extra complexity and parameter overhead associated with prompt tuning and word embeddings required by PriViLege. For miniImageNet, our approach attains stellar performance across all sessions, setting a new benchmark PD of 1.29%, a new base session accuracy of 98.13% and final session accuracy of 96.84% when using the ViT backbone fine-tuned on ImageNet-1K. Interestingly, the less powerful backbone (with ImageNet-1K fine-tuning) outperforms its counterpart, likely due to better alignment with the miniImageNet distribution. Overall, these results validate that our FSCIL framework, with its targeted additive updates and refined classifier weights, effectively maintains robust representations while accommodating new classes, leading to significant improvements over existing methods.